OpenStudy (love_to_love_you):

PLEASE HELP WILL FAN AND MEDAL Write each polynomial function in standard form, classify it by degree, and determine the end behavior of its graph. Show all work. a. y=12-x^4 b. y=2x^3-6x+3x^2-x^4+12 c. y=10-3x^3+3x^2+x^4 They're parts a, b, c, not answers

OpenStudy (love_to_love_you):

@ganeshie8

OpenStudy (micahm):

is this from a test

OpenStudy (love_to_love_you):

no @micahm

ganeshie8 (ganeshie8):

\[\large y=12-x^4\] Write the terms in decreasing exponents for standard form : \[\large y=-x^4+12\]

OpenStudy (anonymous):

what point r u tring to make

ganeshie8 (ganeshie8):

do you know how to find degree of polynomial ?

OpenStudy (love_to_love_you):

Hold on I think I got it

OpenStudy (love_to_love_you):

-x^4+12 quartic binomial down and down ??

ganeshie8 (ganeshie8):

Perfect!

OpenStudy (love_to_love_you):

Yay! I think I got it then! Can you check the other two real quick?

OpenStudy (love_to_love_you):

b. y=-x^4+2x^3+3x^2-6+12; down and down; quartic polynomial

ganeshie8 (ganeshie8):

Yes you got it!

OpenStudy (love_to_love_you):

c. y=x^4-3x^3+3x^2+10; up and up; quartic polynomial

ganeshie8 (ganeshie8):

\[\huge \color{red}{\checkmark}\]

OpenStudy (love_to_love_you):

Great! Thanks! Would you mind helping me with another?

ganeshie8 (ganeshie8):

wil try ask..

OpenStudy (love_to_love_you):

A polynomial function P(x) has degree n. If n is even, is the number of turning points of the graph of P(x) even or odd? What can you say about the number of turning points if n is odd?